Make a study about how many sheets of paper you consumed weekly in


answering your Self Learning Modules. Record the quantity (total number of sheets)


per subject, then construct a probability distribution. Compute the mean, variance,


and the standard deviation of the probability distribution you made. Interpret the


result, then find out how many weeks you will consume 50 sheets of pad paper.


1
Expert's answer
2021-05-07T09:11:01-0400

Solution:

Assume we have the following 4 subjects and we spent the following no. of sheets in them.



Now, mean"=\Sigma X.P(X)=6.44"

And variance"=E(X^2)-[E(X)]^2=42.76-(6.44)^2=1.2864"

Standard deviation"=\sqrt{1.2864}=1.1342"

From the above table, no. of sheets used in a week is 25.

So, 50 sheets will be used in 2 weeks.